A firm has the following revenue and cost functions.

TR = 120 Q – Q2

TC =  1/2Q2 +30 Q + 10

Determine the quantity level at which the firm maximizes its total profit.

(Hint: use marginal revenue = marginal cost rule)       


1
Expert's answer
2021-08-24T17:02:07-0400

"TR=120Q-Q^2"

"MR=120-2Q"


"TC=\frac {1}{2Q^2}+30Q+10"

"MC=\frac {-1}{Q}+30"

At the level at which the firm maximizes profit,

"MR=MC"

"120-2Q=\frac{-1}{Q}+30"

"2Q^2-90Q-1=0"

Solving quadratic equation above:

"Q=\frac{90+/-(8108)^\frac{1}{2}}{2\times2}"

"Q=45"

The quantity level Q=45